### v3

@@ Abstract
This paper provides some foundations for valid forecasting of rare and extreme heat waves through a better understanding of the similarities and differences between several consecutive hot days under normal circumstances and rare, extreme heat waves. We analyze AIRS data from the American Pacific Northwest and AIRS data from the Phoenix, Arizona region. A genetic algorithm is used to help determine the most promising predictors. Classification accuracy with supervised learning is excellent for the Pacific Northwest and is replicated for Phoenix. Conformal prediction sets are considered as a way to represent forecasting uncertainty. Complications caused by endogenous sampling are discussed.
@@ 1 Introduction
In this manuscript, the primary goal is help built a foundation on which valid forecasts of rare and extreme heat waves can be made. To that end, we use a within-subject, quasi-experimental design combined with algorithmic methods in an effort to better distinguish between ordinary, hot summer days and rare, extreme heat waves, perhaps induced by climate change. The data used are a Level 3 product of the NASA Atmospheric Infrared Sounder (AIRS) instrument aboard the Aqua spacecraft (Aumann et al., 2003; Pagano et al., 2010). Observations from the June 2021 heat wave in the North American, Pacific Northwest (PNW) provide an initial forecasting target. Observations from the 2023 July heat wave in the greater Phoenix, Arizona area allow for a replication. For both sites, “control group” comparisons constructed from “usual” summer days are taken from other months or other years for which no credible claims of extreme temperature episodes have been made.
@@
Section 2 addresses some methodological challenges raised by efforts to forecast extreme heat waves, typically characterized by temperatures that likely fall in the far right tail of temperature distributions; they represent rare events that commonly cause serious data sparsity problems. Section 3 provides a grounded discussion of the remotely sensed, AIRS grid cell data to be used and in Section 4, the research designs that can work well with rare events like extreme heat waves are introduced. Substantial progress is made on data sparsity and on design-based adjustments to mitigate the impact of nuisance variables. Section 5 uses gain scores to examine the marginal, ground-level temperature distributions of an actual heat wave compared to a faux heat wave. The distributions are rather different. Section 6 examines whether the predictor distribution differ as well. They do. Section 7 introduces a genetic algorithm to construct a synthetic population of ideal-type, extreme heat wave grid cells, and a synthetic population of ideal-type, ordinary high temperature summer grid cells. Comparisons between the predictor values in these different populations enable a simple form of feature engineering for variable selection. Section 8 uses the selected predictors to examine the promise of forecasting a binary response variable denoting whether or not a grid cell experiences an extreme heat wave. Section 9 reports an effort to reproduce the Pacific Northwest results using an extreme heat wave in the greater Phoenix area. As one would expect, the results somewhat differ, but the overall comparability of the broad conclusions is encouraging. Section 10 returns to the structure of our research designs and the matter of endogenous sampling. We consider the degree to which our efforts to overcome the rare event problem have introduced difficulties in the reported results. We suggest that there is promise in a reformulation from minimizing loss to minimizing risk, but subject-matter and policy considerations are fundamental. Section 11 offers a brief summary of the results and some overall conclusions. The findings are framed as highly provisional in part because many more heat waves in a variety of locations need to be studied.
@@ 2 Some Methodological Challenges
Current global climate models (GCMs) do not appear to offer much help. For the spatial scales at which global climate models function, heat waves are generally below the spatial scales that can be well resolved. Various forms of downscaling seem necessary to construct regional climate models (RCMs), but these introduce a number of new challenges (Gebrechorkos et al., 2023; Lafferty and Sriver, 2023; Schneider et al., 2023). As Gettelman and Rood (2016) write, “For these smaller scale processes, we often must use statistical treatments to match observations to functions that can be used to describe the behavior. Some processes can be represented by basic laws [of physics], other processes must (at the scale of a climate model) be represented by statistical relationships that are only as good as our observations.” To this we add: and only as good as the statistical models themselves (Breiman, 2001b; Freedman, 2009). Moreover, capitalizing on “basic laws” depends on knowing the relevant physics in particular, complex settings, getting boundary conditions from a GCM that are approximately correct, and having access to significant computational resources. And even then, the spatial scale may be too coarse. The analyses to follow suggest that certain quasi-experimental research designs can help when used together with supervised learning.
@@ 4 Research Design
Any research design meant to anticipate extreme heat waves must address the question “compared to what?” The obvious answer is compared to “non-heat waves.” We begin with the 2021 Pacific Northwest extreme heat wave that affected parts of the U.S. and Canada. During the last four days of June, the Pacific Northwest experienced what many observers called an extreme heat wave. Although there are debates over what exactly constitutes an extreme heat wave, this event seems to qualify from almost any informed perspective. With so many extant definitions (Perkins, 2015, Barriopedro et al., 2023), and nothing close to a consensus on best practices, we will provisionally operationalize an extreme heat wave as a single discrete, unusually high temperature event over several days under a slowly moving “heat dome” (Li et al., 2024). More details are provided shortly. Other formulations certainly are possible, and perhaps even preferable, depending on the policy and/or research questions asked.333 The many possible definitions of a heat wave leads to some clumsy wording. The term “true” heat wave implies broad definitional agreement. The same applies to “actual” heat wave. The use of “extreme” heat wave can be quite vague unless “extreme” is defined. Even the term “heat wave” also can mean that we are using a widely accepted definition. To reduce stale repetition, we use all four terms at various times depending on context with the understanding that the definition just provided is what we intend to convey.
@@ 6.2 Some Predictor Comparisons Between the Actual Heat Wave and the Faux Heat Wave 1 or Faux Heat Wave 2
Figure 4 compares predictor histograms from the actual heat wave and Faux Heat Wave 1. At the top is a comparison between distributions for tropopause height plotted red for the true heat wave, in blue for the Faux Heat Wave 1, and in purple for the overlap. The distributions differ in location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, tropopause height was typically substantially higher in the actual heat wave cells. In the middle, there are also differences in distributional location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, atmospheric temperature at level 8 was typically substantially higher in the actual heat wave cells. At the bottom, there are again distributional differences in location, variability and shape. 14 days before both the actual heat wave or Faux Heat wave 1, atmospheric water vapor at level 8 was typically substantially higher in the actual heat wave cells.
One repeated the exercise but replaced Faux Heat Wave 1 with Faux Heat Wave 2. The faux heat wave now is 30 days after the true heat wave. Consequently, the risks from carry-over contamination are dramatically reduced. On the average, however, weather conditions in June and July may differ. There can be less rain and lower soil moisture in July, for instance. According to NOAA National Centers of Environmental Information, 1.9 inches of rain fell in Seattle in June of 2021, and 0.0 inches of rain fell in Seattle in July of 2021 (https://www.ngdc.noaa.gov/).777 In the winter months, the rainfall in Seattle is much more substantial.
@@
There probably are many other variables having distribution differences depending on the circumstances in which summer temperatures materialize. Nevertheless, the comparisons between histograms offer some additional signs that extreme summer heat waves might differ structurally from ordinary summer heat. Work on quasiresonant amplification is a provocative illustration (Petoukov et al., 2012; Mann et al., 2018; Li et al., 2024). However, the empirical details are hard to pin down because there is substantial dependence between many such predictors, especially the when same variable is measured at different altitudes. Proximate values can be highly correlated. It is also important to not uncritically accept any answer to the “compared to what” question. There are no doubt different kinds of “non-heat-waves” to which comparisons can be made. We picked comparisons by temporal and spatial proximity. There are perhaps better, or at least complementary, ways to proceed.
@@ 7 Applying Genetic Algorithms To The Gain Scores
The statistics literature has long suggested, inconclusively, ways to think about and analyze variables that are highly correlated (Alin, 2010). Some recent work seems especially relevant for our setting: variable selection combined with machine learning (Chan et al., 2022). But all approaches we have seen to date rely on the data on hand as the venue for progress. The recommendations are a long way from closure.
@@
We borrow from the social sciences the concept of an “idea type” or a “pure type” (Garth and Mills, 1958) that abstracts the essential features of some phenomenon much like many key concepts in climate science. Then, perhaps a genetic algorithm can be employed to construct two ideal types: a synthetic population of extreme heat wave grid cells and a synthetic population of normal warm summer day grid cells. Comparisons between the two might allow for a more definitive reading of their similarities and differences.
@@ 7.1 Constructing Algorithmic Survival Functions
Random forests handles strongly correlated predictors relatively well because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity.999 Random forests usually randomly samples cases when each tree is grown and predictors at each split in a tree. Random sampling of cases can be done with or without replacement. Random sampling predictors and cases is designed to increase the independence between trees. Greater independence facilitates the averaging over all trees in the forest. We recently learned that gbm3 in R now allows sampling predictors too, and along its R predecessor gbm also allows random sampling of cases. In R, the choice between random forests (i.e., randomForests) gradient boosting now depends largely on details that are peripheral to this discussion. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.101010 We briefly considered using deep neural networks through the keras3 package in R authored and maintained by Tomasz Kalinowski. Unlike random forests algorithm, neither the predictors nor the observations are randomly sampled. The vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, tuning was able to proceed. For all practical purposes, the fitting performance was about as good as random forests’. But, some of the auxiliary procedures available for random forests, which shine some light into the algorithmic black box (Molnar, 2022), were not readily available. For example, the widely applicable partial dependence plots procedure pdp in R, authored and maintained by Brandon Greenwell, apparently requires that a custom prediction function be coded when used on neural networks output.
@@ 7.2.1 Genetic Algorithm Results
The grid cell mean gain score for the final synthetic population produced by the actual heat wave survival function was 7.26 degrees K. The grid cell mean gain score for the final synthetic population produced by the faux heat wave 1 survival function was 5.76 degrees K. Because of differences in the two survival functions, the genetic algorithm did not equate the two. This is another vehicle through which to reflect on whether the physics for a true heat wave may differ somewhat from the physics of a faux heat wave.
@@ 8 Extreme Heat Wave Classification Using a Binary Outcome
Finally, as an intermediate step in the direction of real forecasting, it should be possible for predictors located substantially earlier to identify accurately real heat waves compared to faux heat waves. But accurate classification does not guarantee accurate forecasting; for classification the outcome is known whereas when forecasting, the outcome is not known. (If it were known, there would be no need for a forecast.) New statistical complications can arise.
@@ 8.1 Research Design Revisions
But the rationale goes a bit deeper. When the data are stacked, each study unit no longer serves as its own control. What remains, however, is that the composition of the study units does not change. The same set of grid cells are exposed precursors of a true heat wave and precursors of the faux hear waves. The grid cells serve as the their own collective controls. The confounding would be far more challenging if one set of grid cells were exposed to true heat wave precursors and another set of grid cells were exposed to faux heat wave precursors.
We use the last 4 days of July for the faux heat wave because the time offset between the actual and fax heat wave is 30 days. The chance of carry-over effects becomes very small. This helps justify the absence of a pre-test. Because the earlier results for Faux Heat Wave 1 and Faux Heat Wave 2 were so similar, the different seasonal conditions between June and July probably do not matter much for our classification analysis.
@@ 8.2 Classification Using Random Forests
The earlier comparisons between predictor distributions for a subsequent true heat wave or a subsequent faux heat wave coupled with the genetric algorithm output promise excellent classification results. We considered two forms of widely used and successful supervised learning classifiers: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
@@
Random forests was trained as a classifier using the new data structure. All of the predictors used previously were candidates for the classification exercise. As a start, however, we selected three predictors from Table 2 that were dynamic and seemed most promising: (1) height at the tropopause, (2) atmospheric water vapor mass mixing ratio at level 8, and (3) the temperature at level 8. For each, we considered the gaps shown in Table 2 between values in the synthetic heat wave column and values in the synthetic faux heat wave column. For the mixture and temperature variables, other proximate altitudes could have been selected because proximate altitudes tend to be substantially correlated. Dependence between the three selected predictors by themselves was not a terribly serious complication but still risks some unstable results. We emphasize again that an algorithm is not a model and even if it were, that model would be badly misspecified. There is also the matter of cherry picking the predictors.
@@
The bottom partial dependence plot shows the relationship between heat wave probabilities and the atmospheric water vapor mass mixing ratio at altitude 8. Consistent with its weaker importance to the fit, the relationship shown by the smoother seems haphazard. The fitted values appear disorderly. However, the disappointing performance does not seem to undermine the overall classification accuracy when all three variables are used at once.

### v4

@@ Abstract
This paper provides some foundations for valid forecasting of rare and extreme heat waves through a better understanding of the similarities and differences between several consecutive hot days under normal circumstances and rare, extreme heat waves. We analyze AIRS data from the North American, Pacific Northwest and AIRS data from the Phoenix, Arizona region. A genetic algorithm is used to help determine the most promising predictors. Classification accuracy with supervised learning is excellent for the Pacific Northwest and is replicated for Phoenix. Conformal prediction sets are considered as a way to represent forecasting uncertainty. Complications caused by endogenous sampling are discussed.
@@ 1 Introduction
In this manuscript, the primary goal is to help built a foundation on which valid forecasts of rare and extreme heat waves can be made. The data are taken from the Level 3 product of NASA’s Atmospheric Infrared Sounder (AIRS) instrument aboard the Aqua spacecraft (Aumann et al., 2003; Pagano et al., 2010). Grid cells are the study units. Observations from the June 2021 heat wave in the North American, Pacific Northwest (PNW) are the initial focus. Observations from the 2023 July heat wave in the greater Phoenix, Arizona area allow for a replication. For both sites, “control group” comparisons constructed from “usual” summer days are taken from other months or other years for which no credible claims of extreme temperature episodes have been made.
We offer several contributions.
  1. 1.
    We assemble daily, remote sensing data over many months that include an observed extreme heat wave. In that setting, heat waves are rare events. We employ a within-grid-cell research design that when implemented for sets of chosen days, constructively enables the analysis of rare events.
  2. 2.
    We use as a response variable gain scores of grid cell, surface temperatures to address perhaps the most salient heat wave concern. The gain scores hold constant temporally stable features differing across grid cells (e.g., grid cell elevation), which necessarily cannot be directly associated with the dynamic nature of heat waves. Gains score distributions for the heat wave setting and the control settings are compared. In contrast to much past work, no “off the shelf” statistical distributions are assumed.
  3. 3.
    We also compare the distributions of precursors two weeks before the actual heat wave to distributions of precursors two weeks before ordinary summer days, serving as control settings. Differences between the distributions can be important prerequisites for accurate heat wave forecasts.
  4. 4.
    We then apply a genetic algorithm separately to heat wave data and comparison data to construct two synthetic populations, one for an ideal type extreme wave and one for ideal type ordinary summer days. For each, supervised learning applied to the numeric gain scores provides the requisite fitness functions. Several predictors surface strongly in their ability to differentiate the two synthetic populations and consequently possess forecasting promise, consistent with structural differences between extreme heat waves and commonplace hot summer days. The promise of those predictors is replicated when the response variable is one of two discrete categories: a subsequent heat wave or no subsequent heat wave.
  5. 5.
    Absent any data exploration, we seek to reproduce our results in the greater Phoenix area in order to evaluate the robustness of our research design and substantive findings. The two climate settings could hardly be more different, but the results are qualitatively rather similar.
@@
Section 2 addresses some methodological challenges raised by efforts to forecast extreme heat waves, typically characterized by temperatures that likely fall in the far right tail of temperature distributions; they represent rare events that commonly cause serious data sparsity problems. Section 3 provides a grounded discussion of the remotely sensed, AIRS, grid cell data to be used and in Section 4, the research designs that can work well with rare events like extreme heat waves are introduced. Substantial progress is made on data sparsity and on design-based adjustments to mitigate the impact of nuisance variables. Section 5 uses gain scores to examine the marginal, ground-level temperature distributions of an actual heat wave compared to a faux heat wave. The distributions are rather different. Section 6 examines whether the predictor distribution differ as well. They do. Section 7 introduces a genetic algorithm to construct a synthetic population of ideal-type, extreme heat wave grid cells, and a synthetic population of ideal-type, ordinary high temperature summer grid cells. Comparisons between the predictor values in these different populations enable a simple form of feature engineering for variable selection. Section 8 uses the selected predictors to examine the promise of forecasting a binary response variable denoting whether or not a grid cell experiences an extreme heat wave. Section 9 reports an effort to reproduce the Pacific Northwest results using an extreme heat wave in the greater Phoenix area. As one would expect, the results somewhat differ, but the overall comparability of the broad conclusions is encouraging. Section 10 returns to the structure of our research designs and the matter of endogenous sampling. We consider the degree to which our efforts to overcome the rare event problem have introduced difficulties in the reported results. We suggest that there is promise in a reformulation from minimizing loss to minimizing risk making subject-matter and policy considerations are fundamental. Section 11 offers a brief summary of the results and some overall conclusions. The findings are framed as highly provisional in part because many more heat waves in a variety of locations need to be studied.
@@ 2 Some Methodological Challenges
Current global climate models (GCMs) do not appear to offer much help. For the spatial scales at which global climate models function, heat waves are generally below the spatial scales that can be well resolved. Various forms of downscaling seem necessary in the form of regional climate models (RCMs), but these introduce a number of new challenges (Gebrechorkos et al., 2023; Lafferty and Sriver, 2023; Schneider et al., 2023). As Gettelman and Rood (2016) write, “For these smaller scale processes, we often must use statistical treatments to match observations to functions that can be used to describe the behavior. Some processes can be represented by basic laws [of physics], other processes must (at the scale of a climate model) be represented by statistical relationships that are only as good as our observations.” To this we add: and only as good as the statistical models themselves (Breiman, 2001b; Freedman, 2009). Moreover, capitalizing on “basic laws” depends on knowing the relevant physics in particular, complex settings, getting boundary conditions from a GCM that are approximately correct, and having access to significant computational resources. And even then, the spatial scale may be too coarse. The analyses to follow suggest that certain quasi-experimental research designs can help when used together with supervised learning.
@@ 4 Research Design
Any research design meant to anticipate extreme heat waves must address the question “compared to what?” The obvious answer is compared to “non-heat waves.” We begin with the 2021 Pacific Northwest extreme heat wave that affected parts of the U.S. and Canada. During the last four days of June, the Pacific Northwest experienced what many observers called an extreme heat wave. Although there are debates over what exactly constitutes an extreme heat wave, this event seems to qualify from almost any informed perspective. With so many extant definitions (Perkins, 2015, Barriopedro et al., 2023), and nothing close to a consensus on best practices, we will provisionally operationalize an extreme heat wave as a single discrete, unusually high temperature event over several days under a slowly moving “heat dome” (Li et al., 2024). More details are provided shortly. Other formulations certainly are possible, and perhaps even preferable, depending on the policy and/or research questions asked.333 The many possible definitions of a heat wave leads to some clumsy wording. The term “true” heat wave implies broad definitional agreement. The same applies to “actual” heat wave. The use of “extreme” heat wave can be quite vague unless “extreme” is defined. Even the term “heat wave” also can mean that we are using a widely accepted definition. To reduce stale repetition, we use all four terms at various times depending on context with the understanding that the operationalization just provided is what we intend to convey.
@@ 6.2 Some Predictor Comparisons Between the Actual Heat Wave and the Faux Heat Wave 1 or Faux Heat Wave 2
Figure 4 compares predictor histograms from the actual heat wave and Faux Heat Wave 1. At the top, is a comparison between distributions for tropopause height plotted red for the true heat wave, in blue for the Faux Heat Wave 1, and in purple for the overlap. The distributions differ in location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, tropopause height was typically substantially higher in the actual heat wave grid cells. In the middle, there are also differences in distributional location, variability and shape. 14 days before both the actual heat wave or Faux heat wave 1, atmospheric temperature at level 8 was typically substantially higher in the actual heat wave grid cells. At the bottom, there are again distributional differences in location, variability and shape. 14 days before both the actual heat wave or Faux Heat wave 1, atmospheric water vapor at level 8 was typically substantially higher in the actual heat wave grid cells.
We repeated the exercise but replaced Faux Heat Wave 1 with Faux Heat Wave 2. The faux heat wave now is 30 days after the true heat wave. Consequently, the risks from carry-over contamination are dramatically reduced. On the average, however, weather conditions in June and July may differ. There can be less rain and lower soil moisture in July, for instance. According to NOAA National Centers of Environmental Information, 1.9 inches of rain fell in Seattle in June of 2021, and 0.0 inches of rain fell in Seattle in July of 2021 (https://www.ngdc.noaa.gov/).777 In the winter months, the rainfall in Seattle is much more substantial.
@@
There probably are many other variables having distribution differences depending on the circumstances in which summer temperatures materialize. Nevertheless, the comparisons between histograms offer some additional signs that extreme summer heat waves might differ structurally from ordinary summer heat. Work on quasiresonant amplification is a provocative illustration (Petoukov et al., 2012; Mann et al., 2018; Li et al., 2024).
However, the empirical details are hard to pin down because there is substantial dependence between many such predictors, especially the when same variable is measured at different altitudes. Proximate values can be highly correlated. It is also important to not uncritically accept any answer to the “compared to what” question. There are no doubt different kinds of “non-heat-waves” to which comparisons can be made. We picked comparisons by temporal and spatial proximity. There are perhaps better, or at least complementary, ways to proceed.
@@ 7 Applying Genetic Algorithms To The Gain Scores
The statistics literature has long suggested, inconclusively, ways to think about and analyze variables that are highly correlated (Alin, 2010). Some recent work seems especially relevant for our setting: variable selection combined with machine learning (Chan et al., 2022). But all approaches we have seen to date exploit the data on hand as the venue for progress. The recommendations are a long way from closure.
@@
We borrow from the social sciences the concept of an “idea type” or a “pure type” (Garth and Mills, 1958) that abstracts the essential features of some phenomenon much like many key concepts in climate science. Then, perhaps a genetic algorithm can be employed to construct two ideal types: a synthetic population of extreme heat wave grid cells and a synthetic population of ordinary summer day grid cells. Comparisons between the two might allow for a more definitive reading of their similarities and differences.
@@ 7.1 Constructing Algorithmic Survival Functions
Random forests handles strongly correlated predictors relatively well because it is comprised of a large ensemble of regression or classification trees with the partitioning for each tree undertaken using a small random subsets of predictors. Sampling predictors (without replacement) for each split can dramatically reduce the impact of high multicollinearity.999 Random forests usually randomly samples cases when each tree is grown and predictors at each split in a tree. Random sampling of cases can be done with or without replacement. Random sampling predictors and cases is designed to increase the independence between trees. Greater independence facilitates the averaging over all trees in the forest. We recently learned that gbm3 in R, which implements various kids of boosting, now allows sampling predictors too, and along with its R predecessor gbm allows random sampling of cases as well. In R, the choice between random forests (i.e., randomForests) gradient boosting now depends largely on details that are peripheral to this discussion. We settled on the implementation of random forests in R, ported by Andy Law and Matthew Wiener (Liaw and Weiner, 2002) from the original Fortran by Breiman and Cutler (Breiman, 2001a), which includes several important auxiliary algorithms that can dramatically increase algorithmic transparency.101010 We briefly considered using deep neural networks through the keras3 package in R authored and maintained by Tomasz Kalinowski. Unlike random forests algorithm, neither the predictors nor the observations are randomly sampled. The vastly different units for some of the predictors caused problems for the back propagation. Dependencies between the predictors also caused computational difficulties. Once both problems were addressed, tuning was able to proceed. For all practical purposes, the fitting performance was about as good as random forests’. But, some of the auxiliary procedures available for random forests, which shine some light into the algorithmic black box (Molnar, 2022), were not readily available. For example, the widely applicable partial dependence plots procedure pdp in R, authored and maintained by Brandon Greenwell, apparently requires that a custom prediction function be coded when used on neural networks output.
@@ 7.2.1 Genetic Algorithm Results
The grid cell mean gain score for the final synthetic population produced by the actual heat wave survival function was 7.26 degrees K. The grid cell mean gain score for the final synthetic population produced by the faux heat wave 1 survival function was 5.76 degrees K. Because of differences in the two survival functions, the genetic algorithm did not equate the two. This is another vehicle through which to reflect on how the physics for a true heat wave may differ fundamentally from the physics of a faux heat wave.
@@ 8 Extreme Heat Wave Classification Using a Binary Outcome
Finally, as an intermediate step in the direction of real forecasting, it should be possible for predictors positioned substantially earlier to identify accurately real heat waves compared to faux heat waves. But accurate classification does not guarantee accurate forecasting; for classification the outcome is known whereas when forecasting, the outcome is not known. (If it were known, there would be no need for a forecast.) New statistical complications can arise.
@@ 8.1 Research Design Revisions
But the rationale goes a bit deeper. When the data are stacked, each study unit no longer serves as its own control. What remains, however, is that the composition of the study units does not change. The same set of grid cells are exposed to precursors of a true heat wave and to precursors of the faux hear waves. The grid cells serve as the their own collective controls. The confounding would be far more challenging if one set of grid cells were exposed to true heat wave precursors and another set of grid cells were exposed to faux heat wave precursors.
We use the last 4 days of July for the faux heat wave because the time offset between the actual and faux heat wave is 30 days. The chance of carry-over effects becomes very small. This helps justify the absence of a pre-test. Because the earlier results for Faux Heat Wave 1 and Faux Heat Wave 2 were so similar, the different seasonal conditions between June and July probably do not matter much for our classification analysis.
@@ 8.2 Classification Using Random Forests
The earlier comparisons between predictor distributions for a subsequent true heat wave or a subsequent faux heat wave, coupled with the genetic algorithm output, promise excellent classification results. We considered two forms of widely used and successful supervised learning classifiers: random forests (Breiman, 2001a) and stochastic gradient boosting (Friedman, 2001). Both can be seen as variants of nonparametric regression, are very easy to train, and often the default tuning parameter settings work well. A complication is that there is strong dependence between many of the AIRS predictors. For example, the measured temperatures at different altitudes are often correlated at values above .95.
@@
Random forests was trained as a classifier using the new data structure. All of the predictors used previously were candidates for the classification exercise. As a start, however, we selected three predictors from Table 2 that were dynamic and seemed most promising: (1) height of the tropopause, (2) atmospheric water vapor mass mixing ratio at level 8, and (3) the temperature at level 8. For each, we considered the gaps shown in Table 2 between values in the synthetic heat wave column and values in the synthetic faux heat wave column. For the mixture and temperature variables, other proximate altitudes could have been selected because proximate altitudes tend to be substantially correlated. Dependence between the three selected predictors by themselves was not a terribly serious complication but still risks some unstable results. We emphasize again that an algorithm is not a model and even if it were, that model would be badly misspecified. There is also the matter of cherry picking the predictors.
@@
The bottom partial dependence plot shows the relationship between heat wave probabilities and the atmospheric water vapor mass mixing ratio at altitude 8. Consistent with its weaker importance to the fit, the relationship shown by the smoother seems unsystematic. The fitted values appear disorderly. However, the disappointing performance does not seem to undermine the overall classification accuracy when all three variables are used at once.